# GCLS2: Towards Efficient Community Detection using Graph Contrastive Learning with Structure Semantics

### v1

@@ Abstract
Due to powerful ability to learn representations from unlabeled graphs, graph contrastive learning (GCL) has shown excellent performance in community detection tasks. Existing GCL-based methods on the community detection usually focused on learning attribute representations of individual nodes, which, however, ignores structure semantics of communities (e.g., nodes in the same community should be close to each other). Therefore, in this paper, we will consider the semantics of community structures for the community detection, and propose an effective framework of graph contrastive learning under structure semantics (GCLS2) for detecting communities. To seamlessly integrate interior dense and exterior sparse characteristics of communities with our contrastive learning strategy, we employ classic community structures to extract high-level structural views and design a structure semantic expression module to augment the original structural feature representation. Moreover, we formulate the structure contrastive loss to optimize the feature representation of nodes, which can better capture the topology of communities. Extensive experiments have been conducted on various real-world graph datasets and confirmed that GCLS2 outperforms eight state-of-the-art methods, in terms of the accuracy and modularity of the detected communities.
@@ I Introduction
The community detection (CD) is one of fundamental and important problems in graph data analysis, and it plays an important role in many real-world applications such as the protein function prediction [1], social network analysis [2, 3], anomaly detection [4], and many others. Although traditional CD methods (e.g., hierarchical clustering [5], spectral clustering [6], optimization-based methods [7], etc.) can achieve high accuracy, they often incur high time costs and/or have poor scalability issue for conducting CD over large-scale data graphs, which limits these methods from being widely used in practical applications.
@@
Recently, with the success of unsupervised contrastive learning in image and text representation learning, the graph contrastive learning (GCL) paradigms (e.g., InfoNCE [10], GRACE [11], GCA [12], etc.) have been widely adopted on graphs in various graph-related tasks (e.g., link prediction and node classification), due to their excellent learning capability of the node information. However, these GCL-based methods do not consider structural relationships among nodes in communities, which incurs a reduction in the accuracy and modularity of the detected communities.
@@
Motivation: Figure 1(a) shows a traditional GCL-based (GRACE [11]) community detection framework with two communities (i.e., red and blue groups), which first obtains negative views via data augmentation methods (e.g., edge perturbation, attribute mask, etc.), then constructs positive pairs (e.g., different views of the same nodes) and negative pairs (e.g., different views of neighbor nodes), and finally trains a model to minimize vector distances between positive pairs, while maximizing vector distances between negative pairs. However, intuitively, the feature vector distances of nodes in the red community (same as the blue community) should be close to each other, while the traditional GCL pulls them farther apart. Figure 1(b) shows the heatmaps of the node similarity matrix from a 2-layer DNN applied to the data graph in Figure 1(a), comparing to the results with (left) and without (right) traditional GCL. From Figure 1(b), we observe that nodes expected to belong to the same community (e.g., those within blue and gold boxes) become less similar after GCL training, which will lead to low accuracy and modularity in community detection. This is because the nodes in the same community should be close to each other, while those from different communities should be distant, rather than all nodes in the data graph being pulled apart.
@@
To tackle the aforementioned community detection problem, in this paper, we propose a graph contrastive learning framework under structure semantics for the community detection, named GCLS2. Specifically, we start from the community structure, and apply structure semantic contrastive learning to obtain better feature representation of nodes by extracting structure semantics of graph data. We construct a graph preprocessing module, using existing community dense structures (e.g., $k$-core [13], $k$-truss [14], and $k$-clique [15]) to obtain high-level and original structure views. For different views, we use a structure similarity semantic expression model to obtain structure features of the data graph and design structure contrastive learning to fine-tune node structure feature representations of graph data specific to the community detection task.
@@
  • We propose a Graph Contrastive Learning with Structure Semantics (GCLS2) framework that can effectively extract structural representations of data graphs for community detection.
@@
  • Through extensive experiments, we confirm that our proposed GCLS2 outperforms the supervised and unsupervised learning baselines for community detection.
II Related Work
Graph Contrastive Learning: The graph contrastive learning optimizes the representation of nodes, by maximizing the agreement between pairs of positive samples and minimizing the agreement between pairs of negative samples to be adapted to various downstream tasks. For the node level, there are two general contrastive ways: i) node-graph, for example, DGI [20] contrasts node feature embeddings of augmented view and original data graph with graph feature embeddings of the original data graph, and MVGRL [21] contrasts node feature embeddings of one view with the graph feature embeddings of another view, ii) node-node, for example, GRACE [11] contrasts node feature embeddings of two augmented views with each other, and NCLA [22] contrasts neighbors’ feature embeddings of multiple augmented views. Although these contrastive learning methods have made greater progress in the feature representation, most contrasted samples in the community detection task are against the community’s inherent information representation, where feature embedding representations within communities should be similar and feature embedding representations between communities should be dissimilar. Our proposed GCLS2 approach uses high-level structure adjacency matrix as a signal to guide the anchor closer to dense intra-community and away from the inter-community, and achieves good detection results even when using community structure information only.
III Preliminaries
In this section, we give the definitions of the graph data model and the community detection problem.
III-A Graph Data Model
@@ III-B Community Detection
(Community Detection, CD) Given a data Graph $G$ and a positive integer parameter $L$, the community detection (CD) problem obtains a set of disjoint subgraphs (i.e., communities), $g=\{g_{1},g_{2},\cdots,g_{L}\}$, in the data graph $G$, where each community $g_{i}$ ($1\leq i\leq L$) in $g$ contains classic community structures (e.g., $k$-core, $k$-truss, $k$-plex, etc.).
@@
IV Methodology
@@
IV-A Framework
Figure 2 shows the overall framework of GCLS2. First, we use the classical community structure to extract the high-level structure graph $G^{H}$ from the data graph $G$ for preprocessing, then we obtain the structure similarity semantic vectors of the data graph and the high-level structure graph through structural similarity matrix computation and semantic encoding, and we obtain the feature representations of the two graphs by GCN encoder. After that, we use the feature representations for the structure contrastive learning so that the features better match the structure semantic information of community detection.
IV-B Graph Preprocessing
In GCLS2 model, the graph preprocessing is used to obtain the high-level graph that deeply expresses the structure information of the graph. We show an example of the graph preprocessing in Figure 3. For a data graph $G=(V(G),E(G),\Phi(G),A,X)$, we have the adjacency matrix $A\in\{0,1\}^{N\times N}$ and arrtibute matrix $X\in\mathbb{R}^{N\times F}$, where $N$ denotes the size of $V(G)$ and $F$ denotes the number of attributes. We consider using a set, $M=\{M_{1},M_{2},\cdots,M_{m}\}$, of the classical community structure patterns (i.e., $k$-core, $k$-truss, $k$-plex, etc.) to extract the high-level structure of $G$. Specifically, we count the substructures of the original graph using each pattern $M_{i}\in M$, and we initialize a dictionary $Dic$ to store the number of specific substructural patterns held for each edge pair $\left\langle v_{i},v_{j}\right\rangle$. In Figure 3, we use two substructures (triangle and pentagon, which are 1-core and 3-plex structures, respectively) for pattern counting, and the dashed lines in Figure 3 denote the edges with a pattern count of 0 and masked. In this way, we can obtain a high-level structure graph $G^{H}$ and a high-level adjacency matrix $A^{H}$. It is worth noting that we did not remove those nodes whose neighboring edges all have a pattern number of 0, and retained their attribute values for subsequent structure contrastive learning to maintain model smoothing.
@@ IV-B Graph Preprocessing
IV-C Structure Similarity Semantic Encoder
@@
$S^{\prime}=f(S)=W^{(2)}_{dnn}(\sigma(W^{(1)}_{dnn}S+b^{(1)}_{dnn}))+b^{(2)}_{dnn}$ (2)
$X^{\prime}=f(X)=W^{(2)}_{dnn}(\sigma(W^{(1)}_{dnn}X+b^{(1)}_{dnn}))+b^{(2)}_{dnn}$ (3)
@@
$Z=\mathbb{F}(A,AGG(S^{\prime},X^{\prime}))=\{z_{i}|v_{i}\in V(G)\}$ (4)
@@
$z_{i}^{(l+1)}=\sigma(\sum_{v_{j}\in N(v_{i})}\hat{D}^{-\frac{1}{2}}\hat{A}\hat{D}^{-\frac{1}{2}}z_{j}^{(l)}W^{(l)}_{gcn}+b^{(l)}_{gcn})$ (5)
@@
$z_{i}^{(0)}=AGG(s_{i}^{{}^{\prime}},x_{i}^{{}^{\prime}})|s_{i}^{{}^{\prime}}\in S^{\prime},x_{i}^{{}^{\prime}}\in X^{\prime}$ (6)
@@
$Z^{H}=\mathbb{F}(A^{H},AGG(S^{\prime},X^{\prime}))=\{z_{i}|v_{i}\in V(G^{H})\}$ (7)
@@
IV-D Structure Contrastive Learning
@@
$pos(z_{i}^{1})=\sum\limits_{v_{j}\in N(v_{i})^{H}}(e^{sim(z_{i}^{1},z_{j}^{1})/\tau}+e^{sim(z_{i}^{1},z_{j}^{2})/\tau})$ (8)
$neg(z_{i}^{1})$ $=\sum\limits_{v_{j}\notin N(v_{i})^{H}}(e^{sim(z_{i}^{1},z_{j}^{1})/\tau}+e^{sim(z_{i}^{1},z_{j}^{2})/\tau})$ (9)
$\ell(z_{i}^{1})=-log\frac{(pos(z_{i}^{1})+e^{sim(z_{i}^{1},z_{i}^{2})/\tau})/(2|N(v_{i})^{H}|+1)}{neg(z_{i}^{1})+pos(z_{i}^{1})+e^{sim(z_{i}^{1},z_{i}^{2})/\tau}}$ (10)
@@
Since the two views are symmetric, the node structure contrastive loss of the other view $\ell(z_{i}^{2})$ is defined similarly to Eq. 10. Then we minimize the loss mean of the two views to give the overall comparison loss as follows:
$\mathcal{L}=\frac{1}{2N}\sum_{i=1}^{N}[\ell(z_{i}^{1})+\ell(z_{i}^{2})]$ (11)

### v2

@@ Abstract
Due to the power of learning representations from unlabeled graphs, graph contrastive learning (GCL) has shown excellent performance in community detection tasks. Existing GCL-based methods on the community detection usually focused on learning attribute representations of individual nodes, which, however, ignores structural semantics of communities (e.g., nodes in the same community should be structurally cohesive). Therefore, in this paper, we consider the community detection under the community structure semantics and propose an effective framework for graph contrastive learning under structure semantics (GCLS2) to detect communities. To seamlessly integrate interior dense and exterior sparse characteristics of communities with our contrastive learning strategy, we employ classic community structures to extract high-level structural views and design a structure semantic expression module to augment the original structural feature representation. Moreover, we formulate the structure contrastive loss to optimize the feature representation of nodes, which can better capture the topology of communities. To adapt to large-scale networks, we design a high-level graph partitioning (HGP) algorithm that minimizes the community detection loss for GCLS2 online training. It is worth noting that we prove a lower bound on the training of GCLS2 from the perspective of the information theory, explaining why GCLS2 can learn a more accurate representation of the structure. Extensive experiments have been conducted on various real-world graph datasets and confirmed that GCLS2 outperforms nine state-of-the-art methods, in terms of the accuracy, modularity, and efficiency of detecting communities.
@@ I Introduction
The community detection (CD) is one of fundamental and important problems in graph data analysis, and it plays an important role in many real-world applications such as the protein function prediction [1], social network analysis [2, 3], anomaly detection [4], and many others. Although traditional CD methods (e.g., hierarchical clustering [5], spectral clustering [6], optimization-based methods [7]) can achieve high accuracy, they often incur high time costs and/or have poor scalability issue for conducting CD over large-scale data graphs, which limits these methods from being widely used in practical applications.
@@
Recently, with the success of unsupervised contrastive learning in image and text representation learning, the graph contrastive learning (GCL) paradigms (e.g., InfoNCE [10], GRACE [11], GCA [12], etc.) have been widely adopted in various tasks over graphs such as the link prediction and node classification, due to their excellent learning capability of the node information. However, these GCL-based methods do not consider structural relationships among nodes in communities, which incurs a reduction in the accuracy and modularity of the detected communities.
@@
Motivation: Figure 1(a) shows a traditional GCL-based community detection framework (i.e., GRACE [11]) with two communities (i.e., blue and gold groups), which first obtains negative views via data augmentation methods (e.g., edge perturbation, attribute mask, etc.), then constructs positive pairs (e.g., different views of the same nodes) and negative pairs (e.g., different views of neighbor nodes), and finally trains a model to minimize vector distances between positive pairs, while maximizing vector distances between negative pairs. However, intuitively, the feature vector distances of nodes in the blue community (same as the gold community) should be close to each other, while the traditional GCL pulls them farther apart. Figure 1(b) shows the heatmaps of the node similarity matrix from a 2-layer DNN applied to the data graph in Figure 1(a), comparing to the results without (left) and with (right) traditional GCL. Note that the similarity of each pair of nodes is computed based on the dot product of the embedded representations of the nodes. From Figure 1(b), we observe that the nodes expected to belong to the same community (e.g., those within blue and gold boxes) become less similar after GCL training, which will lead to low accuracy and modularity in the community detection. This is because the embedding representations of nodes in the same community should be close to each other, while those from different communities should be distant, rather than having all nodes in the community being scattered in the GCL embedding space.
@@
To tackle the aforementioned community detection problem, in this paper, we propose a graph contrastive learning (GCL) based framework under structure semantics (named GCLS2) for the community detection. Specifically, we start from the community structure and apply structure semantic contrastive learning to obtain better feature node representations by extracting structure semantics of graph data. We construct a graph preprocessing module, using community dense structures (e.g., $k$-core [13], $k$-truss [14], and $k$-clique [15]) to obtain high-level and original structure views. For different views, we use a structure similarity semantic encoder to obtain structure features of the graph and design structure contrastive learning to fine-tune node structure feature representations of graph data, specific for the community detection task.
Note that, to address the issue that the memory size of local machine limits the information propagation in contrastive learning on large-scale graphs, we design an efficient high-level graph partitioning (HGP) algorithm, which minimizes the loss of potential community integrity while ensuring balanced partitions. Furthermore, from the perspective of the information theory, we demonstrate the effectiveness of our proposed contrastive loss during the training.
@@
  • We propose a Graph Contrastive Learning with Structure Semantics (GCLS2) framework that can effectively extract structural representations of data graphs for the community detection.
@@
  • We devise an efficient high-level graph partitioning (HGP) algorithm to effectively enable the training for contrastive learning over large-scale graphs.
  • We provide a formal proof about the effectiveness of the proposed structure contrastive loss via information theory.
  • Through extensive experiments, we confirm that our proposed GCLS2 approach outperforms the supervised and unsupervised learning baselines on the accuracy, modularity, and efficiency of the community detection.
The remainder of this paper is organized as follows. Section II formally defines the community detection problem in graphs. Section III proposes our GCLS2 model and HGP algorithm, and proves the effectiveness of our GCLS2 model training. Section IV conducts comprehensive experiments to compare our GCLS2 approach with eight state-of-the-art algorithms. Section V introduces related works on the community detection and graph contrastive learning. Finally, Section VI concludes this paper.
II Problem Definition
In this section, we give formal definitions of the graph data model and community detection problem.
II-A Graph Data Model
@@ II-B The Community Detection
(Community Detection, CD) Given a data Graph $G$ and a positive integer parameter $L$, the community detection (CD) problem obtains a set of disjoint subgraphs (i.e., communities), $g=\{g_{1},g_{2},\cdots,g_{L}\}$, in the data graph $G$, where each community $g_{i}$ ($1\leq i\leq L$) in $g$ contains classic community structures (e.g., $k$-core, $k$-truss, $k$-plex).
@@
III Graph Contrastive Learning with Structure Semantics for the Community Detection
@@
III-A The GCLS2 Framework
Figure 2 shows the overall GCLS2 framework for the community detection over a data graph $G$. We first use classical community structures to extract a high-level structure graph $G^{H}$ from the data graph $G$ for preprocessing, then we obtain the structure similarity semantic vectors of the data graph and the high-level structure graph through structural similarity matrix computation and semantic encoding, and we obtain the feature representations of the two graphs by GCN encoder. After that, we use the feature representations for the structure contrastive learning so that the features better match the structure semantic information of community detection.
III-B Graph Preprocessing
In GCLS2 model, the graph preprocessing is used to obtain the high-level graph that deeply expresses the structure information of the graph. We show an example of the graph preprocessing in Figure 3. For a data graph $G=(V(G),E(G),\Phi(G),A,X)$, we have the adjacency matrix $A\in\{0,1\}^{N\times N}$ and arrtibute matrix $X\in\mathbb{R}^{N\times F}$, where $N$ denotes the size of $V(G)$ and $F$ denotes the number of attributes. We consider using a set, $M=\{M_{1},M_{2},\cdots,M_{m}\}$, of the classical community structure patterns (i.e., $k$-core, $k$-truss, $k$-plex) to extract the high-level structure of $G$. Specifically, we count the substructures of the original graph using each pattern $M_{i}\in M$, and we initialize a dictionary $Dic$ to store the number of specific substructural patterns held for each edge pair $\left\langle v_{i},v_{j}\right\rangle$. In Figure 3, we use two substructures (triangle and pentagon, which are 1-core and 3-plex structures, respectively) for pattern counting, and the dashed lines in Figure 3 denote the edges with a pattern count of 0 and masked. In fact, other specified dense structures (e.g., rectangles) could also be used for the task. However, in this paper, we focus on pattern counting based on the two fundamental structures mentioned above. In this way, we can obtain a high-level structure graph $G^{H}$ and a high-level adjacency matrix $A^{H}$. It is worth noting that we did not remove those nodes whose neighboring edges all have a pattern number of 0, and retained their attribute values for subsequent structure contrastive learning to maintain model smoothing.
@@ III-B Graph Preprocessing
III-C Structure Similarity Semantic Encoder
@@
$S^{\prime}=f(S)=W^{(2)}_{dnn}(\sigma(W^{(1)}_{dnn}S+b^{(1)}_{dnn}))+b^{(2)}_{dnn},$ (2)
$X^{\prime}=f(X)=W^{(2)}_{dnn}(\sigma(W^{(1)}_{dnn}X+b^{(1)}_{dnn}))+b^{(2)}_{dnn},$ (3)
@@
$Z=\mathbb{F}(A,AGG(S^{\prime},X^{\prime}))=\{z_{i}|v_{i}\in V(G)\}.$ (4)
@@
$z_{i}^{(l+1)}=\sigma(\sum_{v_{j}\in N(v_{i})}\hat{D}^{-\frac{1}{2}}\hat{A}\hat{D}^{-\frac{1}{2}}z_{j}^{(l)}W^{(l)}_{gcn}+b^{(l)}_{gcn}),$ (5)
@@
$z_{i}^{(0)}=AGG(s_{i}^{{}^{\prime}},x_{i}^{{}^{\prime}})|s_{i}^{{}^{\prime}}\in S^{\prime},x_{i}^{{}^{\prime}}\in X^{\prime}.$ (6)
@@
$Z^{H}=\mathbb{F}(A^{H},AGG(S^{\prime},X^{\prime}))=\{z_{i}|v_{i}\in V(G^{H})\}.$ (7)
@@
III-D Structure Contrastive Learning
@@
$pos(z_{i}^{1})=\sum\limits_{v_{j}\in N(v_{i})^{H}}(e^{sim(z_{i}^{1},z_{j}^{1})/\tau}+e^{sim(z_{i}^{1},z_{j}^{2})/\tau}),$ (8)
$neg(z_{i}^{1})$ $=\sum\limits_{v_{j}\notin N(v_{i})^{H}}(e^{sim(z_{i}^{1},z_{j}^{1})/\tau}+e^{sim(z_{i}^{1},z_{j}^{2})/\tau}),$ (9)
$\ell(z_{i}^{1})=-log\frac{pos(z_{i}^{1})+e^{sim(z_{i}^{1},z_{i}^{2})/\tau}}{neg(z_{i}^{1})+pos(z_{i}^{1})+e^{sim(z_{i}^{1},z_{i}^{2})/\tau}},$ (10)
@@
Since the two views are symmetric, the node structure contrastive loss of the other view $\ell(z_{i}^{2})$ is defined similarly to Eq. 10. Then we minimize the loss mean of the two views to give the overall contrastive loss as follows:
$\mathcal{L}=\frac{1}{2N}\sum_{i=1}^{N}[\ell(z_{i}^{1})+\ell(z_{i}^{2})].$ (11)
III-E The High-Level Graph Partitioning
For large-scale graphs, the memory of most machines cannot keep up with the increase of the number of neighbors of a GNN, making many graph representation methods like graph contrastive learning difficult to execute on GPUs.
Existing graph partitioning algorithms for large-scale graph learning mainly have two categories, vertex partitioning [22, 23, 24] and edge partitioning [25, 26]. Note that, the edge partitioning assigns edges to the corresponding subgraphs, leading to vertex redundancy and multiple representations of some vertices during the graph learning. This redundancy increases computational overhead for query processing. Thus, in this paper, we instead adopt a vertex partitioning approach, which assigns vertices to subgraphs in large-scale data graphs.
The two objectives of the graph partitioning are load balancing and minimizing the cost of cutting. Optimizing these two objectives together is called the Balanced Graph Partitioning problem, which is NP-hard [27]. Traditional vertex partitioning algorithm, METIS [22], uses the sparsification and refinement aggregation of consecutive vertices and edges to partition the data graph, but requires random traversal and scaling of the entire graph structure, which tends to break the community structure, resulting in a large loss of boundary features. The LGD [23] algorithm uses a greedy strategy to aggregate neighboring nodes in the same partition and ensures load balancing in each partition. As shown in Figure 1(a), not all neighboring nodes belong to the same community. The partitions generated by the LDG algorithm can lead to abnormal community detection results. Moreover, METIS and LDG perform poorly on large-scale data graphs even though they try to keep the load balanced as much as possible.
To adapt to our community detection task, we propose a framework for high-level graph partitioning (HGP) to minimize the damage to the community structure, which is illustrated in Figure 4.
The HGP Framework: Our HGP framework consists of three main steps: (1) high-level graph partitioning, (2) rebuilding of the original data graph, and (3) partition training of the model.
Note that, cutting edges may break the community structure and affect the downstream detection results, but this is unavoidable. Therefore, our goal of HGP is to minimize the loss of the community structure while ensuring load balancing in graph partitioning, defined as follows.
Definition 3.
(High-Level Graph Partitioning, HGP) Given a data graph $G(V,E)$, the number, $p$, of partitions, and the balanced partition rate $a$, the objectives of HGP are given by:
$\begin{cases}\max_{i\in[1,p]}|V(part_{i})|\leq\frac{(1+a)|V(G)|}{p}\\ \min|E(G)-(\bigcup_{i\in[1,p]}E(part_{i}))|\end{cases},$ (12)
where $V(part_{i})$ and $E(part_{i})$ denote the sets of vertices and edges in partition $part_{i}$, respectively.
To achieve this goal, we design an HGP optimization algorithm as shown in Algorithm 1. Specifically, we first obtain a high-level adjacency matrix $A^{H}$ precomputed in Section III-B for high-level structural partitioning of the data graph $G$ (line 1). Next, we make a judgment on each vertex $v_{i}\in V(G)$: if the current vertex $v_{i}$ has no neighbors in the high-level adjacency matrix $A^{H}$, i.e., this is a solitary vertex, then we consider assigning $v_{i}$ to the partition with the highest number of neighbors in the original data graph $G$ (lines 2-4). As shown in the partition data graph $G^{P}$ in Figure 4, this yellow solitary vertex in the upper left corner is assigned to partition $P^{H}_{1}$ because all its original neighbors are in partition $P^{H}_{1}$. Otherwise, we test each partition $j=1,\cdots,p$: if the partitioning size is less than the load limit, we query how many high-level neighbors of $v_{i}$ exist in the current partition $part_{j}$, we aim to divide $v_{i}$ into the partition with the highest number of high-level neighbors, so we compute the current partition score,
$s(v_{i},part_{j})=|part_{j}\bigcap N(v_{i})^{H}|\cdot\frac{|V(G)|}{p(|part_{j}|+1)},$ (13)
where the second term is a load limit to limit the size of the partition (lines 5-9). Finally, we choose to assign $v_{i}$ to the partition with the highest score $s$ (line 10) until we return all partitioning results for each vertex (line 11).
After completing the high-level graph partitioning, we perform the mapping partitioning of the original graph to rebuild the data graph. Specifically, we synchronously map the vertex partition IDs obtained from the HGP algorithm into the original data graph, resulting in $p$ partitions with the same size as those obtained by the HGP algorithm. As shown in Figure 4, the $v_{1}^{H}$ and $v_{2}^{H}$ in $G^{P}$ are assigned to different partitions, and after mapping, the original data graph cuts the edge $e(v_{1},v_{2})$ to form the rebuild data graph $G^{\prime}$.
Actually, we obtain $p$ high-level structural subgraphs and $p$ original subgraphs after the HGP partitioning and rebuild. In the training process of contrastive learning, we take each pair of subgraphs (i.e., $\left\langle P_{1}^{H},P_{1}\right\rangle$) for contrastive training. Since the size of the cut subgraphs can be controlled by the specified parameter $p$, users can completely adjust the parameter $p$ through the local memory for the online contrastive learning training. In Figure 4, $\left\langle P_{1}^{H},P_{1}\right\rangle$ and $\left\langle P_{2}^{H},P_{2}\right\rangle$ are learned by the same GCLS2 model, which results in the corresponding loss values $\ell_{1}$ and $\ell_{2}$. To tune the learning globally, we design the following global loss for backpropagation learning through these loss values.
$\mathcal{L}_{P}=\sum\limits_{p}\frac{|V(P_{p})|}{N}\ell_{p}.$ (14)
It is worth noting that since the training is on different subgraph pairs and our partitioning results maximally guarantee the load balancing of each partition through the HGP algorithm. Therefore, we can further improve the training efficiency of large-scale data graphs by parallel training.
Input: i) the data graph $G$, and ii) the number, $p$, of partitions
Output: graph partitions in $pid$
1
[…]
3for each vertex $v_{i}\in V(G)$ do
4
      // solitary vertex
5
6      if $N(v_{i})^{H}=\emptyset$ then
7
8            $pid[v_{i}]=\mathop{\arg\max}\limits_{j\in\{1,\cdots,p\}}\{|part_{j}\bigcap N(v_{i})|\}$
9      else
10
11            for each partitions $j=1,\cdots,p$ do
12
13                  if $|part_{j}|<\frac{|V(G)|}{p}$ then
14
                        // high-level neighbors in partition $j$
15
16                        $part_{j}\bigcap N(v_{i})^{H}$
                        // partition score with load limit
17
18                        $s(v_{i},part_{j})=|part_{j}\bigcap N(v_{i})^{H}|\cdot\frac{|V(G)|}{p(|part_{j}|+1)}$
19
20
21            $pid[v_{i}]=\mathop{\arg\max}\limits_{j\in\{1,\cdots,p\}}\{s(v_{i},part_{j})\}$
22
23
return $pid$
Algorithm 1 HGP Algorithm
III-F Model Training and Complexity Analysis
The Model Training: The detailed GCLS2 training process is shown in Algorithm 2. Specifically, we first obtain the high-level structure graph $G^{H}$ and high-level adjacency matrix $A^{H}$ from the original data graph $G$, and construct a dictionary $Dic$ by counting the number of specified patterns for each edge of $G$, then extract the structure similarity matrix $S$ from the $Dic$ (lines 1-3). For each training iteration, we first use the SSS module to extract structural semantics representation $S^{\prime}$ and attribute semantics representation $X^{\prime}$ (lines 4-6). Next, we employ GCNs with weight shared to extract node features from both $G$ and $G^{H}$, resulting in the node feature matrix $Z$ and the high-level node feature matrix $Z^{H}$ (lines 7-8). Then, we calculate the similarity between positive samples and the similarity between negative samples to obtain the structure contrastive loss $(\ell(z_{i}^{1}),\ell(z_{i}^{2}))$ for each vertex $v_{i}$ (line 9). After that, we update the GCLS2 network parameters by minimizing all structure contrastive loss $\mathcal{L}$ on $Z$ and $Z^{H}$ (lines 10-11). Finally, after $\varepsilon$ training iterations, we return the node features $Z$ obtained for subsequent community detection (line 12).
The Complexity Analysis: We analyze the time complexity and space complexity of the GCLS2 model by the number of floating point operations (FLOPs) and space overhead required by the GCLS2 model. Specifically, let the time required for each unit floating point operation be $O(1)$. For the Structure Similarity Semantic Encoder module, the time complexity of the SSS module is $O(Nd^{2})$, and the time complexity of the node representation extractor is $O(|E|d)$. For the Structure Contrastive Learning module, the similarity of each pair of vertices can be realized by vector dot product, which would cost $O(d)$. Then, the time complexity of all structure contrastive loss is $O(Nd)$. Therefore, the time complexity of the training process of GCLS2 is $O(Nd^{2}+|E|d)$
For the Structure Similarity Semantic Encoder module, the space complexity is $O(N^{2}+d^{2})$, which contains the input graph size and the SSS parameter size. Similarly, the space complexity of the node representation extractor is $O(Nd+d^{2})$. For the Structure Contrastive Learning module, the computation of the structure contrastive loss can be performed on the feature matrices $Z$ and $Z^{H}$ with a space complexity of $O(N^{2})$. Since the individual training process is serial and $d$ is significantly smaller than $N$, the space complexity of the training process of GCLS2 is $O(N^{2})$.
For the HGP algorithm, the time complexity of the partitioning is $O(pN+|E|)$. The space complexity of the partitioning is $O(N)$,
1